# Enhancing LLM-based Test Generation for Hard-to-Cover Branches via Program Analysis

### v1

@@ IV-A3 Variants for RQ2
To answer RQ2, we constructed five variants of TELPA for the ablation experiment:
TELPA that removes the backward method-invocation analysis from TELPA;
TELPA that removes the forward method-invocation analysis from TELPA;
@@
Similar to RQ1, we constructed two sets of variants based on the preceding test generation tool: Pynguin-based variants (TELPA, TELPA, TELPA, TELPA, TELPA); and CODAMOSA-based variants (TELPA, TELPA, TELPA, TELPA, TELPA).
IV-A4 Variants for RQ3
@@
TELPA is built on top of the adopted preceding test generation tools and LLM. To better understand how this design choice will impact the performance of TELPA, in this RQ, we construct another two variants. In the first variant, named TELPA, no preceding test generation tool is used. That is, TELPA is activated from the very beginning. In the second variant, we consider a different LLM. The default LLM in TELPA is Phind-CodeLlama-34B-v2, which is relatively large-scale and may require a lot of resources to run. To make TELPA more practical, we also investigated the effectiveness of TELPA when a relatively small-scale LLM is adopted. More specifically, we employed DeepSeek-Coder-6.7B-instruct [21], which is one of the most effective LLMs among the set of LLMs with comparable (small) scales on HuggingFace. Accordingly, we call the variant TELPA. Similar to previous RQs, Pynguin and CODAMOSA based variants are created and named as TELPA and TELPA.
IV-A5 Benchmark, Metrics, and Environment
@@
Following the existing work [4], we evaluated TELPA on the widely-studied benchmark, which consists of 486 modules from 27 Python projects. The whole benchmark contains 42897 source lines of code and 4518 methods in total.
Following the existing work on test generation [1, 4], we adopted both branch coverage and line coverage as metrics to measure the effectiveness of a test generation technique. For each module and each technique in a project, we allocated 20 minutes for test generation. All the techniques kept running until the time budget ran out, even if the coverage did not grow before that.
@@
To avoid the influence of randomness, we repeated our experiments for three times considering the evaluation cost and used the average value of the final results. The experiments were conducted on a workstation with 128-core CPU, 504G memory, 4 NVIDIA A100 GPUs, and Ubuntu 20.04 OS.
IV-B RQ1: Effectiveness of TELPA
| Project | Pynguin | Pynguin | BasicPrompt | TELPA | CODAMOSA | CODAMOSA | BasicPrompt | TELPA |
| pysnooper | 16.95% | 18.64% | 21.47% | 29.94% | 17.07% | 17.68% | 23.00% | 31.10% |
| apimd | 39.88% | 43.30% | 55.76% | 81.31% | 7.14% | 40.06% | 9.43% | 67.39% |
| blib2to3 | 23.68% | 26.24% | 24.02% | 42.59% | 19.75% | 25.72% | 20.47% | 38.59% |
| codetiming | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% |
| cookiecutter | 46.28% | 55.37% | 63.64% | 65.29% | 45.45% | 57.02% | 52.16% | 66.12% |
| dataclasses_json | 16.55% | 16.55% | 28.42% | 28.06% | 17.27% | 17.63% | 29.50% | 28.42% |
| docstring_parser | 43.66% | 53.52% | 54.23% | 85.21% | 84.65% | 85.82% | 90.07% | 92.20% |
| flutes | 67.46% | 75.40% | 70.63% | 84.92% | 66.94% | 75.81% | 67.63% | 81.45% |
| flutils | 46.29% | 48.00% | 68.86% | 78.57% | 66.01% | 66.29% | 68.91% | 78.37% |
| httpie | 28.03% | 29.69% | 46.74% | 45.59% | 26.42% | 27.20% | 44.64% | 44.64% |
| isort | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% |
| mimesis | 77.16% | 80.50% | 79.94% | 89.69% | 68.83% | 79.11% | 72.73% | 89.42% |
| py_backwards | 28.44% | 41.74% | 36.70% | 63.76% | 30.77% | 35.75% | 36.65% | 60.63% |
| pymonet | 64.44% | 65.93% | 76.30% | 85.93% | 61.07% | 64.12% | 76.34% | 88.55% |
| pypara | 43.75% | 56.82% | 46.02% | 63.64% | 18.02% | 48.84% | 20.14% | 36.63% |
| semantic_release | 38.64% | 38.64% | 45.45% | 60.23% | 36.78% | 36.78% | 49.43% | 58.62% |
| string_utils | 83.93% | 85.71% | 93.45% | 98.21% | 64.52% | 80.00% | 72.28% | 87.88% |
| pytutils | 36.11% | 36.67% | 41.67% | 57.78% | 40.68% | 44.63% | 45.76% | 57.06% |
| sanic | 43.96% | 45.33% | 56.59% | 62.64% | 46.17% | 48.36% | 54.10% | 60.11% |
| sty | 89.13% | 91.30% | 93.48% | 95.65% | 91.30% | 94.48% | 91.30% | 95.65% |
| thefuck | 19.64% | 20.36% | 31.79% | 50.71% | 22.01% | 23.66% | 33.16% | 53.76% |
| thonny | 17.95% | 36.32% | 32.91% | 37.61% | 26.75% | 36.75% | 33.16% | 42.31% |
| tornado | 46.33% | 47.97% | 58.10% | 65.19% | 42.95% | 44.60% | 53.24% | 65.44% |
| tqdm | 14.62% | 19.23% | 19.23% | 44.62% | 35.29% | 41.18% | 37.18% | 50.00% |
| typesystem | 32.69% | 37.59% | 57.21% | 51.11% | 20.81% | 55.84% | 24.12% | 72.59% |
| youtube_dl | 15.34% | 17.03% | 15.34% | 24.90% | 20.89% | 23.54% | 20.89% | 29.11% |
| ansible | 27.87% | 29.73% | 28.19% | 38.03% | 29.76% | 31.36% | 29.84% | 37.84% |
| Average Branch Cov. | 44.77% | 48.80% | 53.56% | 64.12% | 44.72% | 51.93% | 50.23% | 63.48% |
| Average Line Cov. | 63.23% | 66.02% | 68.85% | 76.36% | 59.13% | 67.83% | 62.55% | 75.51% |
@@ IV-B RQ1: Effectiveness of TELPA
Table I presents the comparison results for TELPA vs. Pynguin vs. BasicPrompt and TELPA vs. CODAMOSA vs. BasicPrompt on each project in terms of branch coverage. To gain a better understanding of the performance of the preceding test generation step, we also show the branch coverage achieved by Pynguin and CODAMOSA right before switching to LLMs in “Pynguin” and “CODAMOSA” columns, respectively. “bn” represents “bottleneck”. The last two rows present the average branch coverage and line coverage across all projects. Due to the space limit, we do not show the line coverage for each project.
@@
To better understand the effectiveness of TELPA and TELPA, we also measured the time spent on running Pynguin and CODAMOSA before switching to LLMs. On average across all projects, Pynguin and CODAMOSA ran for 2.57 and 3.51 minutes, respectively, before switching. That is, TELPA is activated at an early stage of the entire testing process (given a time budget of 20 minutes) regardless of taking Pynguin or CODAMOSA as the preceding test generation tool. When comparing Pynguin, Pynguin, CODAMOSA, and CODAMOSA, we found that on average, Pynguin achieves a branch coverage of 44.77% within only 2.57 minutes, but just achieves 48.80% coverage after continuing running for 17.43 minutes. Similarly, CODAMOSA achieves 44.72% branch coverage within only 3.51 minutes, but in the end just achieves 51.93% coverage after continuing running for 16.49 minutes. That is, Pynguin and CODAMOSA can quickly cover easily reachable branches, and then encounter the coverage bottleneck to some degree. This highlights the challenge posed by hard-to-cover branches and underscores the motivation behind our work.
@@
After switching to TELPA, the branch coverage can be significantly improved according to results of TELPA and TELPA. Specifically, after reaching the testing time budget, TELPA achieves an average improvement of 31.39% over Pynguin across all projects in terms of branch coverage, and TELPA achieves an average improvement of 22.22% over CODAMOSA. In terms of line coverage, TELPA achieves an average improvement of 15.67% over Pynguin and TELPA achieves an average improvement of 11.31% over CODAMOSA. Furthermore, we conducted a paired Wilcoxon signed-rank test [22] at a significance level of 0.05 between TELPA/TELPA and Pynguin/CODAMOSA, confirming TELPA’s statistical superiority over the two state-of-the-art techniques with all p-values below 0.05.
@@
Regardless of the preceding test generation tools (Pynguin or CODAMOSA), TELPA can achieve similar effectiveness, i.e., TELPA achieves similar branch coverage and line coverage to TELPA on average across all projects. The results demonstrate the stable effectiveness of TELPA. We conducted a paired Wilcoxon signed-rank test [22] at a significance level of 0.05 between TELPA and TELPA in terms of both branch coverage and line coverage. Both p-values exceed 0.05, demonstrating no statistically significant difference between TELPA and TELPA in improving coverage.
@@
Compared to Pynguin, all the LLM-based techniques (including TELPA, TELPA, BasicPrompt, BasicPrompt, CODAMOSA) achieve higher branch coverage and line coverage on average across all projects. This demonstrates the effectiveness of LLMs in improving test coverage. However, different prompting methods can largely affect the effectiveness of LLMs. By comparing TELPA/TELPA with BasicPrompt/BasicPrompt, our designed prompting method specific to the challenge posed by hard-to-cover branches performs much better than the general prompting method employed for LLM-based test generation (e.g., CODAMOSA). Specifically, the average improvement of TELPA over BasicPrompt is 19.71% in terms of branch coverage and 10.92% in terms of line coverage. Similarly, the average improvement of TELPA over BasicPrompt is 26.38% in terms of branch coverage and 20.71% in terms of line coverage. This implies the importance of designing task-specific prompting, i.e., extracting useful information (rather than leveraging as much available information as possible) for improving the coverage of hard-to-cover branches in TELPA. We conducted a paired Wilcoxon signed-rank test [22] at a significance level of 0.05 and found that TELPA/TELPA significantly outperforms BasicPrompt/BasicPrompt by obtaining all p-values smaller than 0.05.
IV-C RQ2: Contribution of Each Component in TELPA
| TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | |
| Branch Cov. | 56.32% | 58.82% | 55.44% | 58.72% | 58.60% | 64.12% | 50.79% | 53.66% | 50.31% | 51.15% | 51.58% | 63.48% |
| Line Cov. | 70.61% | 72.20% | 70.34% | 71.60% | 71.51% | 76.36% | 62.69% | 64.46% | 62.83% | 63.05% | 63.32% | 75.51% |

### v2

@@ 4.1.3. Variants for RQ2
To answer RQ2, we constructed seven variants of TELPA for the ablation experiment to investigate the contribution of each component in TELPA :
TELPA that removes the object construction analysis from TELPA;
TELPA that removes the branch dependency analysis from TELPA;
@@
TELPA that removes the path filtering process from TELPA. Specifically, instead of condensing the sequences with the same entry and ending method, we keep all sequences.
TELPA that removes the functionality summarization process from TELPA.
@@
Similar to RQ1, we constructed two sets of variants based on the preceding test generation tool: Pynguin-based variants (TELPA, TELPA, TELPA, TELPA, TELPA, TELPA, TELPA); and CODAMOSA-based variants (TELPA, TELPA, TELPA, TELPA, TELPA, TELPA, TELPA).
4.1.4. Variants for RQ3
@@
TELPA is built on top of the adopted preceding test generation tools and LLM. To better understand how this design choice will impact the performance of TELPA, in this RQ, we construct another two variants. In the first variant, named TELPA, no preceding test generation tool is used. That is, TELPA is activated from the very beginning. In the second variant, we consider a different LLM. The default LLM in TELPA is Phind-CodeLlama-34B-v2, which is relatively large-scale and may require a lot of resources to run. To make TELPA more practical, we also investigated the effectiveness of TELPA when a relatively small-scale LLM is adopted. More specifically, we employed DeepSeek-Coder-6.7B-instruct (dee, 2024), which is one of the most effective LLMs among the set of LLMs with comparable (small) scales on HuggingFace according to the LLM leaderboard maintained by Hugging Face (lea, 2024). Accordingly, we call the variant TELPA. Similar to previous RQs, Pynguin-based and CODAMOSA-based variants are created and named as TELPA and TELPA. Furthermore, the commercial closed-source LLMs developed by OpenAI, such as ChatGPT and GPT-4, have shown superior performance on various benchmarks. Therefore, in addition to open-source LLMs, we also conducted an experiment to understand the effectiveness of TELPA when using the more advanced GPT-4. Due to the high cost of invoking GPT-4’s APIs, we only constructed a Pynguin-based variant, i.e., TELPA.
4.1.5. Baseline for RQ5
To answer RQ5, we adapted TELPA to Java and constructed TELPAe, which leverages the state-of-the-art search-based test generation tool for Java (i.e., EvoSuite (Fraser and Arcuri, 2011b)) for preceding test generation and compared it with EvoSuite.
EvoSuite is one of the most prominent search-based techniques. It employs evolutionary algorithms to generate unit tests aimed at maximizing code coverage. Additionally, it integrates advanced program analysis techniques to further enhance its effectiveness, including leveraging a constant pool for generating test inputs and applying testability transformations to improve guidance during the test generation process, etc.
4.1.6. Benchmark, Metrics, and Environment
@@
To answer RQ1-RQ4, following the existing work (Lemieux et al., 2023), we evaluated TELPA on the widely-studied benchmark, which consists of 486 modules from 27 Python projects. The whole benchmark contains 42,897 source lines of code and 4,518 methods in total. Since TELPA focuses on improving coverage for hard-to-cover branches, we also analyzed the branches within the benchmark. The benchmark contains a total of 6,559 branches, of which 3,979 are classified as hard-to-cover. Specifically, 3,931 branches involve complex dependencies, while 256 branches are associated with complex objects.
To answer RQ5, we further evaluated TELPA on four real-world open-source projects using the JUnit framework (jun, 2024) in the widely-used Defects4J benchmark (Just et al., 2014) due to the significant popularity of Java and JUnit (i.e., Chart, Time, Lang, and Math). Unlike Python, Java is a statically typed language. Automated test case generation for dynamically typed languages often struggles with object creation, as obtaining detailed type information can be challenging. In contrast, Java’s static typing allows for more precise and efficient handling of type information, making it an ideal choice for evaluating the effectiveness of TELPA on static type systems. Here, we excluded Closure since it does not use the JUnit framework and involves significant testing costs. In total, the benchmark contains 9,391 branches, of which 5,494 are classified as hard-to-cover. Specifically, 2,453 branches involve complex dependencies, while 4,842 branches are associated with complex objects.
Following the existing work on test generation (Lukasczyk and Fraser, 2022; Lemieux et al., 2023), we adopted both branch coverage and line coverage as metrics to measure the effectiveness of a test generation technique. Specifically, to evaluate the correctness of the generated tests cases in RQ4, we executed all generated test cases and measure the correctness using two metrics: (1) Syntax Correctness refers to the percentage of generated tests that are syntactically valid. Specifically, we checked this by parsing the generated tests using Python’s built-in ast module. Any test that passed parsing without errors was considered syntactically correct. (2) Execution Pass Rate refers to the percentage of generated tests that executed successfully without runtime errors. Specifically, we ran the tests and recorded any runtime errors or failures during execution. For each module and each technique in a project, we allocated 20 minutes for test generation. All the techniques kept running until the time budget ran out, even if the coverage did not grow before that. Note that the time costs spent on all components in TELPA (including the preceding test generation, program analysis, test generation, and the coverage analysis) were included in this time budget. The 20-minute time budget was selected to ensure that all the studied techniques have sufficient time to explore the test space and demonstrate their potential as much as possible. While the previous studies set shorter time budgets for evaluation (e.g., 10 minutes for evaluating Pynguin and CODAMOSA), a longer time budget is helpful to assess the sustained effectiveness of each technique over a sufficient period.
@@
To avoid the influence of randomness, we repeated our experiments for 20 times considering the evaluation cost and used the average value of the final results. The experiments were conducted on a workstation with 128-core CPU, 504G memory, 4 NVIDIA A100 GPUs, and Ubuntu 20.04 OS.
We released our implementation and all experimental data at the project homepage (hom, 2025) to facilitate replication, future research, and practical use.
4.2. RQ1: Effectiveness of TELPA
Project | Pynguin | Pynguin | CHATTESTER | TELPAp | CODAMOSA | CODAMOSA | CHATTESTER | TELPAc |
pysnooper | 16.67% | 18.07% | 20.00% | 31.23% | 16.67% | 18.25% | 27.37% | 30.70% |
apimd | 39.63% | 43.18% | 48.04% | 81.50% | 6.92% | 39.81% | 51.78% | 67.94% |
blib2to3 | 23.55% | 26.09% | 25.18% | 42.39% | 19.59% | 25.58% | 30.66% | 38.78% |
codetiming | 65.00% | 75.00% | 80.00% | 95.00% | 70.00% | 70.00% | 60.00% | 95.00% |
cookiecutter | 45.37% | 54.15% | 63.90% | 64.63% | 45.37% | 57.80% | 52.44% | 67.07% |
dataclasses_json | 16.11% | 16.37% | 21.77% | 27.96% | 17.17% | 17.61% | 17.17% | 28.94% |
docstring_parser | 43.58% | 53.77% | 72.64% | 85.85% | 83.77% | 86.04% | 89.62% | 91.13% |
flutes | 68.61% | 77.22% | 81.94% | 84.17% | 68.06% | 76.67% | 76.11% | 82.50% |
flutils | 46.52% | 47.41% | 59.70% | 78.89% | 66.22% | 66.74% | 67.26% | 78.00% |
httpie | 28.14% | 29.79% | 38.35% | 45.52% | 26.49% | 27.11% | 40.05% | 44.90% |
isort | 94.00% | 99.00% | 98.00% | 100.00% | 94.00% | 98.00% | 94.00% | 98.00% |
mimesis | 76.97% | 80.42% | 80.25% | 89.75% | 68.57% | 79.58% | 83.45% | 89.66% |
py_backwards | 28.62% | 42.00% | 50.00% | 63.69% | 31.69% | 36.00% | 33.23% | 60.31% |
pymonet | 62.67% | 65.33% | 68.00% | 85.33% | 61.33% | 63.33% | 71.56% | 88.22% |
pypara | 44.13% | 56.35% | 47.14% | 63.97% | 18.41% | 48.57% | 34.13% | 56.51% |
semantic_release | 37.50% | 38.33% | 45.56% | 60.28% | 37.22% | 37.22% | 39.72% | 60.00% |
string_utils | 83.75% | 85.62% | 93.12% | 97.97% | 64.84% | 80.62% | 85.31% | 87.97% |
pytutils | 35.95% | 37.03% | 50.95% | 57.57% | 41.08% | 45.14% | 43.92% | 57.57% |
sanic | 44.56% | 44.91% | 59.65% | 63.07% | 46.23% | 48.60% | 51.58% | 59.91% |
sty | 87.14% | 90.00% | 91.43% | 95.71% | 90.00% | 93.57% | 89.29% | 95.00% |
thefuck | 19.88% | 20.71% | 21.07% | 50.60% | 21.90% | 23.45% | 38.69% | 54.17% |
thonny | 17.35% | 36.63% | 34.58% | 38.31% | 27.11% | 36.51% | 37.95% | 43.13% |
tornado | 46.30% | 47.89% | 55.75% | 65.36% | 42.79% | 44.05% | 51.20% | 65.61% |
tqdm | 15.00% | 18.48% | 40.45% | 44.70% | 35.30% | 42.12% | 38.33% | 49.24% |
typesystem | 33.08% | 37.50% | 39.46% | 50.87% | 20.94% | 56.09% | 57.32% | 72.54% |
youtube_dl | 15.28% | 16.97% | 19.46% | 25.03% | 20.70% | 23.57% | 24.56% | 29.01% |
ansible | 28.00% | 29.86% | 30.88% | 38.02% | 29.73% | 31.38% | 30.76% | 37.78% |
Average Branch Cov. | 43.09% | 47.71% | 53.23% | 63.98% | 43.41% | 50.87% | 52.50% | 64.06% |
Average Line Cov. | 62.36% | 64.72% | 70.20% | 75.54% | 58.37% | 67.92% | 72.64% | 74.25% |
Branch Type | Pynguin | Pynguin | CHATTESTER | TELPAp | CODAMOSA | CODAMOSA | CHATTESTER | TELPAc | |
Easy-to-cover | 64.50% | 72.51% | 77.17% | 78.84% | 73.10% | 78.18% | 80.65% | 82.32% | |
Hard-to-cover | Branches with complex dependencies | 29.00% | 31.16% | 32.15% | 39.10% | 18.75% | 20.17% | 28.80% | 34.90% |
Branches with complex objects | 52.73% | 57.03% | 58.98% | 63.67% | 39.06% | 42.97% | 53.90% | 58.59% | |
All hard-to-cover branches (union of the above two types) | 29.20% | 31.62% | 32.67% | 39.53% | 19.13% | 20.58% | 29.10% | 35.23% |
@@ 4.2. RQ1: Effectiveness of TELPA
The goal of TELPA is to reach hard-to-cover branches, and this RQ verifies its ability in this regard. Table 1 presents the comparison results for TELPAp vs. Pynguin vs. CHATTESTER and TELPAc vs. CODAMOSA vs. CHATTESTER on each project in terms of branch coverage. To gain a better understanding of the performance of the preceding test generation step, we also show the branch coverage achieved by Pynguin and CODAMOSA right before switching to LLMs in “Pynguin” and “CODAMOSA” columns333Here, “bn” represents “bottleneck”., respectively. The last two rows present the average branch coverage and line coverage across all projects. Due to the space limit, we do not show the line coverage for each project but instead include these detailed results in our replication package (hom, 2025). To provide a clearer comparison of those techniques on hard-to-cover branches, Table 2 specifically shows the branch coverage achieved by each technique on different types of branches.
@@
To better understand the effectiveness of TELPAp and TELPAc, we also measured the time spent on running Pynguin and CODAMOSA before switching to LLMs. On average across all projects, Pynguin and CODAMOSA ran for 2.39 and 3.47 minutes, respectively, before switching. That is, TELPA is activated at an early stage of the entire testing process (given a time budget of 20 minutes) regardless of taking Pynguin or CODAMOSA as the preceding test generation tool. When comparing Pynguin, Pynguin, CODAMOSA, and CODAMOSA, from Table 1, we found that on average, Pynguin achieves a branch coverage of 43.09% within only 2.39 minutes, but just achieves 47.71% coverage after continuing running for 17.61 minutes. Similarly, CODAMOSA achieves 43.41% branch coverage within only 3.47 minutes, but in the end just achieves 50.87% coverage after continuing running for 16.53 minutes. From Table 2, this trend is consistent for hard-to-cover branches as well, where Pynguin and CODAMOSA quickly encounter bottlenecks in achieving further coverage improvements on hard-to-cover branches within a short time period. Figure 8 shows the branch-coverage trend achieved by TELPAp, Pynguin, TELPAc, and CODAMOSA. From this figure, Pynguin and CODAMOSA can quickly cover easily reachable branches, and then encounter the coverage bottleneck to some degree, while TELPAp and TELPAc continue to achieve new coverage. This highlights the challenge posed by hard-to-cover branches and underscores the motivation behind our work. This also confirms that the two-minute timeframe set for switching to LLMs is sufficient for Pynguin and CODAMOSA, which can reflect their potential in covering hard-to-cover branches to a large extent, as extending their running time does not bring significant coverage improvement. Additionally, over the 20-minute period, TELPA’s coverage continues to increase, though at a slower rate after the first 10 minutes, underscoring the importance of sufficient testing time to fully harness the testing capabilities. In contrast, Pynguin and CODAMOSA quickly plateau within the first few minutes, while TELPA steadily achieves higher coverage. Notably, TELPA also surpasses both Pynguin and CODAMOSA within the 10-minute time budget, demonstrating its effectiveness.
@@
From Table 1, after switching to TELPA, the branch coverage can be significantly improved according to results of TELPAp and TELPAc. Specifically, after reaching the testing time budget, TELPAp achieves an average improvement of 34.10% over Pynguin across all projects in terms of branch coverage, and TELPAc achieves an average improvement of 25.93% over CODAMOSA. In terms of line coverage, TELPAp achieves an average improvement of 16.72% over Pynguin and TELPAc achieves an average improvement of 9.32% over CODAMOSA. Additionally, TELPAp outperforms CODAMOSA in terms of both branch and line coverage. Specifically, TELPAp achieves an average branch coverage of 63.98% while CODAMOSA achieves an average branch coverage of 50.87%. This also demonstrates the superiority of the prompting method and the test generation process designed in TELPA. Regarding hard-to-cover branches, as shown in Table 2, both TELPAp and TELPAc outperform Pynguin and CODAMOSA on both scenarios, including those with complex dependencies and those with complex objects. For example, TELPAp covers 39.53% hard-to-cover branches, outperforming Pynguin, which covers 31.62% branches with an improvement of 25.02%. Similarly, TELPAc covers 35.23% hard-to-cover branches, compared to CODAMOSA’s 20.58% branches, achieving an improvement of 71.18%. Furthermore, we conducted a paired Wilcoxon signed-rank test (Woolson, 2007) at a significance level of 0.05 between TELPAp/TELPAc and Pynguin/CODAMOSA, confirming TELPA’s statistical superiority over the two state-of-the-art methods with all p-values below 0.05.
@@
Regardless of the preceding test generation tools (Pynguin or CODAMOSA), TELPA can achieve similar effectiveness, i.e., TELPAp achieves similar branch coverage and line coverage to TELPAc on average across all projects. The results demonstrate the stable effectiveness of TELPA. We conducted a paired Wilcoxon signed-rank test (Woolson, 2007) at a significance level of 0.05 between TELPAp and TELPAc in terms of both branch coverage and line coverage. Both p-values exceed 0.05, demonstrating no statistically significant difference between TELPAp and TELPAc in improving coverage.
@@
Compared to Pynguin, all the LLM-based/LLM-enhanced techniques (including TELPAp, TELPAc, CHATTESTER, CHATTESTER, CODAMOSA) achieve higher branch coverage and line coverage on average across all projects. This demonstrates the effectiveness of LLMs in improving test coverage. However, different prompting methods can largely affect the effectiveness of LLMs. By comparing TELPAp/TELPAc with CHATTESTER/CHATTESTER, our designed prompting method specific to the challenge posed by hard-to-cover branches performs much better than the general prompting method employed for LLM-based test generation. Specifically, the average improvement of TELPAp over CHATTESTER is 20.19% in terms of branch coverage and 7.61% in terms of line coverage. Similarly, the average improvement of TELPAc over CHATTESTER is 22.02% in terms of branch coverage and 2.22% in terms of line coverage. The improvement is also consistent on hard-to-cover branches. Specifically, the improvement of TELPAp over CHATTESTER is 21.00% and the improvement of TELPAc over CHATTESTER is 21.06% on hard-to-cover branches. This implies the importance of designing task-specific prompting, i.e., extracting relevant information for improving the coverage of hard-to-cover branches in TELPA. We conducted a paired Wilcoxon signed-rank test (Woolson, 2007) at a significance level of 0.05 and found that TELPAp/TELPAc significantly outperforms CHATTESTER/CHATTESTER by obtaining all p-values smaller than 0.05.
To gain deeper insights into the specific branches covered by each method, we analyzed the overlap in covered branches across different methods, as visualized in Figure 9. The figure demonstrates that all studied techniques perform well on easy-to-cover branches. However, TELPA not only covers nearly all the hard-to-cover branches covered by other methods but also covers a significant number of unique hard-to-cover branches. This underscores TELPA’s contribution to enhancing coverage of hard-to-cover branches and complementing existing methods.
4.3. RQ2: Contribution of Each Component in TELPA
Cov. | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPAp |
Branch Cov. | 55.87% | 58.74% | 55.17% | 58.34% | 58.69% | 52.80% | 59.84% | 63.98% |
Line Cov. | 70.67% | 71.80% | 70.46% | 71.65% | 71.07% | 69.17% | 72.29% | 75.54% |
Cov. | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPAc |
Branch Cov. | 56.68% | 60.45% | 56.12% | 57.77% | 57.32% | 48.89% | 60.12% | 64.06% |
Line Cov. | 72.52% | 73.78% | 72.24% | 72.30% | 71.89% | 66.88% | 72.48% | 74.25% |